20kg of N2 and 3kg of H2 are mixed to produce a NH3. The amount of NH3 formed is?

Grade:12th pass

7 Answers

Gursimran Singh
8 years ago
It is a compond not a mixture so 50 kg is not the answer ,consider hydrogen= 2 and nitrogen = 14 ,And here are 3 hydrogen so 2 * 3=6, 6:14 = 3:7 if 3 kg is given and our ratio is too 3 kg so 7 kg of n2 will make compound and compound mass would be 10 kg
Gursimran Singh
8 years ago
But simply adding can also give approx. value 17030 + 5994 = 23000 grams approx.Do by conversion in moles and also use the law!
Aamir Nabi bhat
6 years ago
N2+3H2=2NH3NO. Of moles ofn2=20*1000/281=714.2No of moles of H2=30*100/2=150002*714.2=1428.4Mass of NH3=1428.4*17=20Kg
Rahul chouhan
6 years ago
It is a compond not a mixture so 50 kg is not the answer ,consider hydrogen= 2 and nitrogen = 14 ,And here are 3 hydrogen so 2 * 3=6, 6:14 = 3:7 if 3 kg is given and our ratio is too 3 kg so 7 kg of n2 will make compound and compound ...#mass would be 10 kg
AJ
5 years ago
Ans
AJ
5 years ago
Ans is 17 kg. 28g of N2 gives 6g of H2 according to eq. So 3000g of H2 gives (3000*28)/6 = 14000g. Since 28g of N2 makes 34g NH3 therefore 14000g N2 makes 17000g or 17 kg NH3 ie (14000*34)/28
Chandrakant Senad
5 years ago
N2 + 3H2 ― 2NH3
Since, no. of moles= wt. in gram÷ atomic mass
Moles of N2= 20× 10^(-3) ÷ 28
                      = 0.7×10^(-3)
Moles of H2= 3×10^(-3)÷2
                      = 1.5×10^-3
1 mole of N2 reacts with 3 moles of H2
0.7×10^(-3) moles N2 reacts with 2.1× 10^(-3) moles H2. {Moles of H2= 0.7×10^(-3)× 3}
 
This implies that, H2 is L.R. (Limiting Reagent)
 
A/C Chemical Reaction:-
3 moles of H2 forms 2 moles of NH3
1.5 ×10^(-3) moles of H2 forms [2/3×1.5×10^)(-3)]
                                             = 10^(-3) moles of NH3
Therefore, no. Of moles of NH3= 10^(-3)
Since, moles= wt. In gram ÷ atomic mass
      So, wt. In gram = moles × atomic mass
                                  = 17 × 10^(-3) g
            So, Wt. In kg= 17 Kg